# On irrationals with Lagrange value exactly 33

### v1

@@ 1.1. Lagrange Spectrum and bi-infinite sequences
In this paper, $\mathbb{N}=\{0,1,2,\dots\}$ represents the set of non-negative integer numbers and $\mathbb{N}^{*}=\{1,2,\dots\}$ represents the set of positive integer numbers. Notice that through the continued fraction, we can identify $x$ with an infinite sequence $(x_{i})_{i\in\mathbb{N}}$, and we are defining Lagrange value for infinite sequences in (1). We can extend the definition of Lagrange value to bi-infinite sequences: Consider the symbolic dynamical system $(\Sigma,\sigma)$, where $\Sigma=(\mathbb{N}^{*})^{\mathbb{Z}}$ is the symbolic space (also the space of bi-infinite sequences) and $\sigma:\Sigma\rightarrow\Sigma,(x_{i})_{i\in\mathbb{Z}}\mapsto(x_{i-1})_{i\in\mathbb{Z}}$ is the left shift. For any $\underline{x}=(x_{i})_{i\in\mathbb{Z}}\in\Sigma$ , define the height function $\lambda$ on $\Sigma$ as
@@ 1.2. Statement of the main result
These facts follows from very classical results (Malyshev1981MarkovAL and Markoff1879 ,Markoff1880 ), but the case $c=3$ is apparently not present in the literature. This paper is filling that gap: we prove that the set $X_{3}$ is also uncountable.
@@
$\pi:\Sigma=(\mathbb{N}^{*})^{\mathbb{Z}}\rightarrow\Sigma^{+}=(\mathbb{N}^{*})^{\mathbb{N}},\quad\underline{x}=(x_{i})_{i\in\mathbb{Z}}\mapsto\pi(\underline{x})=(x_{i})_{i\in\mathbb{N}}$
@@ 2.1. Markov’s Theorem
Markov’s theorem gives a complete description of the spectrum $\mathcal{L}$ and $\mathcal{M}$ at the part before $3$, here we state a version in Aigner2013 .
@@
$\mathcal{L}\cap(-\infty,3)=\mathcal{M}\cap(-\infty,3)=\left\{\sqrt{5},\sqrt{8},\frac{\sqrt{221}}{5},\dots,\frac{\sqrt{9z^{2}-4}}{z},\dots\right\}$
where $z$ is called Markov number, which is the largest coordinate of positive integer solutions of Markov equation $z_{1}^{2}+z_{2}^{2}+z_{3}^{2}=3z_{1}z_{2}z_{3}$.
[…]
@@
The continued fractions of quadratics are eventually periodic ((see (Aigner2013, , Theorem 1.17) for a proof)). In Markov’s Theorem, the period of $\gamma_{z}$ is given by the Cohn tree ((see Figure 1)), i.e., every period of $\gamma_{z}$ is some vertex of the Cohn tree and every vertex of the Cohn tree is some period of $\gamma_{z}$. Without loss of generality, we can choose the $\gamma_{z}$ in the theorem to be purely periodic.
@@ 2.2. Structure of Admissible Bi-infinite Words
Let $x_{1};x_{2},\dots]$ and $y_{1};y_{2},\dots]$ be two infinite continued fractions, then $x_{1};x_{2},\dots]<[y_{1};y_{2},\dots]$ if and only if $(-1)^{k+1}x_{k}<(-1)^{k+1}y_{k}$, where $k$ is the first index such that $x_{k}\neq y_{k}$.
We use a vertical bar between two letters to represent a cut of word (of any type). For example, if $x=x_{0}x_{1}x_{2}\dots$ is an infinite word, the cut at the $i+1$-th position is $x_{0}\dots x_{i-1}x_{i}|x_{i+1}\dots$. We define the value of a cut of infinite or bi-infinite words by
@@
We can also define the notation of transpose of words. Let $x=x_{0}x_{1}\dots x_{n}$ be a finite word, we define the transpose of $x$ to be $x^{T}=x_{n}x_{n-1}\dots x_{0}$. Similarly, let $x=x_{0}x_{1}\dots$ be an infinite word and $\underline{x}=\dots x_{-1}x_{0}x_{1}\dots$ be a bi-infinite word, their transpose are defined by $x^{T}=\dots x_{1}x_{0}$ (infinite to the left) and $\underline{x}^{T}=\dots x_{1}x_{0}x_{-1}\dots$ (transpose at 0-th position).
@@ 3.1. Ideas of the construction
Since we can identify irrationals with infinite sequences and the definition of Lagrange value for irrationals can be naturally extended to infinite words, we will not distinguish irrationals with infinite words in the following text.
@@
$(\alpha,\beta)$$(\alpha\beta,\beta)$$(\alpha\beta\beta,\beta)$$(\alpha\beta\beta\beta,\beta)$$(\alpha\beta\beta\beta\beta,\beta)$$(\alpha\beta\beta\beta,\alpha\beta\beta\beta\beta)$$(\alpha\beta\beta,\alpha\beta\beta\beta)$$(\alpha\beta,\alpha\beta\beta)$$(\alpha,\alpha\beta)$ $\overline{U}$$\overline{V}$$\overline{U}$$\overline{V}$$\overline{U}$$\overline{V}$$\overline{U}$$\overline{V}$
@@ 3.3. Uncountably many elements in $Y_{3}(0)$
Let $\alpha_{n}^{T},\beta_{n}^{T}$ be factors of some elements in $\pi(m^{-1}(3))$, then all the cuts inside $\alpha_{n}^{T},\beta_{n}^{T}$ are good except the cuts at the last letter $2$.
@@ 3.4. Uncountably many elements in $Y_{3}(n)$ for $n\in\mathbb{N}_{+}$
Clearly the cuts at the first $n$ positions (at the letter $3$) are bad cuts, so we only need to prove that there are no bad cuts inside the suffix $\omega_{X_{3}(0)}$. The proof proceeds by induction, which is similar to Lemma 3.6. We only need to check that every odd cuts inside the suffix $\omega_{X_{3}(0)}$, the rest part of the induction coincide with the proof of Lemma 3.6.
@@
where we are using the fact that $\omega_{0}\in\{1,2\}$ is smaller than $3$ and the $i$-th position is a odd position. ∎
However, it is evident that the elements we found in Proposition 3.8 are not projections of elements in $m^{-1}(3)$. To prove that the set $Y_{3}(n)$ is uncountable, one way is to anticipate a $\alpha_{n}^{T}$ and remove the first letter $b$ at the beginning of $\omega_{X_{3}(0)}$, i.e., consider the word $(\alpha_{n}^{T})^{+}\omega_{X_{3}(0)}$.
@@ 4. Related Questions
It is also interesting to study the structure of $\widetilde{\mathcal{M}}$. The first observation is that, different from the Lagrange value $k$, the function $\widetilde{m}$ depends on the initial part of the continued fraction, so maybe the spectrum will not be as “good” as the Lagrange spectrum or Markov spectrum. The first element of $\widetilde{\mathcal{M}}$ is no longer $\sqrt{5}$, in fact, it is $\widetilde{m}([1;1,1,\dots])=\frac{3+\sqrt{5}}{2}$. But similarly with $\mathcal{L}$ and $\mathcal{M}$, is the initial part of $\widetilde{\mathcal{M}}$ also a discrete sequence accumulating precisely at $3$? Does $\widetilde{\mathcal{M}}$ has positive Hausdorff dimension? Does $\widetilde{\mathcal{M}}$ has the Hall’s ray?

### v2

@@ 1.1. Lagrange Spectrum and bi-infinite sequences
In this paper, $\mathbb{N}=\{0,1,2,\dots\}$ represents the set of non-negative integer numbers and $\mathbb{N}^{*}=\{1,2,\dots\}$ represents the set of positive integer numbers. Notice that through the continued fraction, we can identify $x$ with an infinite sequence $(x_{i})_{i\in\mathbb{N}}$, and we are defining Lagrange value for infinite sequences in (1). We can extend the definition of Lagrange value to bi-infinite sequences: Consider the symbolic dynamical system $(\Sigma,\sigma)$, where $\Sigma=(\mathbb{N}^{*})^{\mathbb{Z}}$ is the symbolic space (also the space of bi-infinite sequences) and $\sigma:\Sigma\rightarrow\Sigma,(x_{i})_{i\in\mathbb{Z}}\mapsto(x_{i+1})_{i\in\mathbb{Z}}$ is the left shift. For any $\underline{x}=(x_{i})_{i\in\mathbb{Z}}\in\Sigma$ , define the height function $\lambda$ on $\Sigma$ as
@@ 1.2. Statement of the main result
These facts follows from very classical results (Malyshev1981MarkovAL and Markoff1879 ,Markoff1880 ), but the case $c=3$ is apparently not present in the literature. In fact, Hurwitz Hurwitz claimed that $X_{3}$ is infinite, and more recently, in (gugu, , Chapter 8) it was claimed that $X_{3}$ is uncountable, however, both references presented these claims without providing proofs. This paper is filling that gap: we prove that the set $X_{3}$ is also uncountable.
@@
$\pi:\Sigma=(\mathbb{N}^{*})^{\mathbb{Z}}\rightarrow\Sigma^{+}=(\mathbb{N}^{*})^{\mathbb{N}^{*}},\quad\underline{x}=(x_{i})_{i\in\mathbb{Z}}\mapsto\pi(\underline{x})=(x_{i})_{i\in\mathbb{N}^{*}}$
@@ 2.1. Markov’s Theorem
Definition 2.3.
A Markov number $m$ is the largest coordinate of a solution of the Markov equation $x^{2}+y^{2}+z^{2}=3xyz$ in the positive integers. The multiplicity of a Markov number $m$ is the number of distinct triples $(x,y,z)$ such that $x,y,z$ are positive integers, $x^{2}+y^{2}+z^{2}=3xyz$ and $x<=y<=z=m$.
A well known uniqueness conjecture by Frobenius states that all the Markov numbers have multiplicity of $1$ and the conjecture still remains open.
Markov’s theorem gives a complete description of the spectrum $\mathcal{L}$ and $\mathcal{M}$ at the part before $3$. Here we state a version in Bom .
@@
$\mathcal{L}\cap(-\infty,3)=\mathcal{M}\cap(-\infty,3)=\left\{\sqrt{5},\sqrt{8},\frac{\sqrt{221}}{5},\dots,\frac{\sqrt{9m^{2}-4}}{m},\dots\right\}$
where $m$ is a Markov number.
Moreover, let $(m_{r})_{r\geq 1}$ be the sequence of the ordered Markov numbers with multiplicity, i.e., a non-decreasing sequence whose terms are all Markov numbers and such that the number of times each term appears is equal to its multiplicity. There is a sequence of $GL_{2}(\mathbb{Z})$-inequivalent irrationals $\gamma_{r}=\frac{a_{r}+\sqrt{9{m_{r}}^{2}-4}}{b_{r}}$ $(a_{r},b_{r}\in\mathbb{Z})$ such that $k(\gamma_{r})=\frac{\sqrt{9m_{r}^{2}-4}}{m_{r}}$, and every $x\in\mathbb{R}\setminus\mathbb{Q}$ with $k(x)<3$ is $GL_{2}(\mathbb{Z})$-equivalent to some $\gamma_{r}$.
@@
The continued fractions of quadratics are eventually periodic ((see (Bom, , Theorem A.1) for a proof)). In Markov’s Theorem, the period of $\gamma_{r}$ is given by the Cohn tree ((see Figure 1)), i.e., every period of $\gamma_{r}$ is some vertex of the Cohn tree and every vertex of the Cohn tree is some period of $\gamma_{r}$. Without loss of generality, we can choose the $\gamma_{r}$ in the theorem to be purely periodic.
@@ 2.2. Structure of Admissible Bi-infinite Words
Let $x_{0};x_{1},\dots]$ and $y_{0};y_{1},\dots]$ be two infinite continued fractions, then $x_{0};x_{1},\dots]<[y_{0};y_{1},\dots]$ if and only if $(-1)^{k}x_{k}<(-1)^{k}y_{k}$, where $k$ is the first index such that $x_{k}\neq y_{k}$.
We use a vertical bar between two letters to represent a cut of word (of any type). For example, if $x=x_{1}x_{2}x_{3}\dots$ is an infinite word, the cut at the $i+1$-th position is $x_{1}\dots x_{i-1}x_{i}|x_{i+1}\dots$. We define the value of a cut of infinite or bi-infinite words by
@@
We can also define the notation of transpose of words. Let $x=x_{1}x_{2}\dots x_{n}$ be a finite word, we define the transpose of $x$ to be $x^{T}=x_{n}\dots x_{2}x_{1}$. Similarly, let $x=x_{1}x_{2}\dots$ be an infinite word and $\underline{x}=\dots x_{-1}x_{0}x_{1}\dots$ be a bi-infinite word, their transpose are defined by $x^{T}=\dots x_{2}x_{1}$ (infinite to the left) and $\underline{x}^{T}=\dots x_{1}x_{0}x_{-1}\dots$ (transpose at 0-th position).
@@ 3.1. Ideas of the construction
Given $x=x_{1}x_{2}\dots$ an infinite word, we can always identify $x$ with the irrational $[0;x]$, and we can naturally extend the definition of Lagrange value for irrationals to infinite words, so we will not distinguish between infinite word $x$ and irrational $[0;x]$ in the following text.
@@
$(\alpha,\beta)$$(\alpha\beta,\beta)$$(\alpha\beta\beta,\beta)$$(\alpha\beta\beta\beta,\beta)$$(\alpha\beta\beta\beta\beta,\beta)$$(\alpha\beta\beta\beta,\alpha\beta\beta\beta\beta)$$(\alpha\beta\beta,\alpha\beta\beta\beta)$$(\alpha\beta,\alpha\beta\beta)$$(\alpha,\alpha\beta)$ $\tilde{U}$$\tilde{V}$$\overline{U}$$\overline{V}$$\overline{U}$$\overline{V}$$\overline{U}$$\overline{V}$$\overline{U}$$\overline{V}$
@@ 3.3. Uncountably many elements in $Y_{3}(0)$
Let $\alpha_{n}^{T},\beta_{n}^{T}$ be factors of some elements in $\pi(m^{-1}(3))$ occurring at an odd position, then all the cuts inside $\alpha_{n}^{T},\beta_{n}^{T}$ are good except the cuts at the last letter $2$.
@@ 3.4. Uncountably many elements in $Y_{3}(n)$ for $n\in\mathbb{N}_{+}$
Clearly the cuts at the first $n$ positions (at the letter $3$) are bad cuts, so we only need to prove that there are no bad cuts inside the suffix $\omega_{X_{3}(0)}$. The proof proceeds by induction, which is similar to Lemma 3.6. We only need to check that every odd cut inside the suffix $\omega_{X_{3}(0)}$ is good, the rest part of the induction coincide with the proof of Lemma 3.6.
@@
where we are using the fact that $\omega_{0}\in\{1,2\}$ is smaller than $3$ and the $i$-th position is an odd position. ∎
We have now completed the proof of Theorem 1.2. However, it is evident that the elements we found in Proposition 3.8 are not projections of elements in $m^{-1}(3)$. To prove that the set $Y_{3}(n)$ is uncountable, one way is to anticipate an $\alpha_{n}^{T}$ and remove the first letter $b$ at the beginning of $\omega_{X_{3}(0)}$, i.e., consider the word $(\alpha_{n}^{T})^{+}\omega_{X_{3}(0)}$.
@@ 4. Related Questions
It is also interesting to study the structure of $\widetilde{\mathcal{M}}$. The first observation is that, different from the Lagrange value $k$, the function $\widetilde{m}$ depends on the initial part of the continued fraction, so maybe the spectrum will not be as “good” as the Lagrange spectrum or Markov spectrum. The first element of $\widetilde{\mathcal{M}}$ is no longer $\sqrt{5}$, in fact, it is $\widetilde{m}([1;1,1,\dots])=\frac{3+\sqrt{5}}{2}$. But similarly with $\mathcal{L}$ and $\mathcal{M}$, is the initial part of $\widetilde{\mathcal{M}}$ also a discrete sequence accumulating precisely at $3$? Does $\widetilde{\mathcal{M}}$ has positive Hausdorff dimension? Does $\widetilde{\mathcal{M}}$ have Hall’s ray?